# Impact of data usage for forecasting on performance of model predictive control in buildings with smart energy storage

### v1

@@ 3.3.3 Online training
During building operation monitoring systems continuously collect operational data. The characteristics of building behaviour can change during operation due to external factors such as weather & climate, occupancy, and equipment degradation & maintenance. Continuous online training updates the predictive models using the collected monitoring data, improving the model’s ability to adapt to dynamic changes in building behaviour. A version of the Linear prediction model with online training is implemented to investigate the impact of online training frequency on model accuracy. The prediction accuracy of models updated at different frequencies is plotted in Fig. 9. Higher update frequencies led to increased prediction accuracy across all prediction variables. For example, prediction accuracy for grid electricity price improved by 15.7%, 14.0%, 12.1%, and 8.1% when updated monthly, quarterly, semi-annually, and annually, respectively, compared to the baseline model which is not trained online. Therefore, it is expected that the additional computational hardware and system complexity required for online trained prediction models will be worthwhile in practical systems.
@@ 3.4 Sensitivity of Model Predictive Control to forecast accuracy
Whether investments in additional data usage to improve forecast accuracy provide net benefit to the operation of a smart energy storage system depends on the resulting improvements in operational performance achieved by the MPC controller. The determination of optimal data usage strategies therefore requires quantification of the relationship between forecast accuracy and MPC operational performance. To investigate this in a controlled setting, the operational performance achieved by the tested MPC scheme is evaluated with the use of synthetic forecasts. The synthetic forecasts are produced by adding a Gaussian random walk noise component to the ground-truth values, as described in Eq. 4,
@@
The results show that the tested MPC scheme is most sensitive to the forecast accuracy of the grid electricity price and carbon intensity variables, suggesting that the most resource and expense should be invested in producing accurate forecasts of the grid conditions. Additionally, whilst the prediction models tested in this study performed worst when forecasting solar generation, see Fig. 2b, it may not be worth expending additional resource to improve these forecasts as the MPC scheme is least sensitive to this variable. Though the grid component of operational performance is found to be by far the most sensitive to forecast accuracy, this is likely due to the synthetic forecast model used [^3] In the synthetic forecast model used, the Gaussian random walk noise component is regenerated at each time instance. This leads to the generation of forecasts which are correct on average, but can cause the MPC controller to take opposing energy management strategies in successive time steps, leading to large magnitudes of energy flows from the battery to correct the strategy. Further work is required to create a noise model for synthetic forecasts that is calibrated to the operational performance vs forecast accuracy trade-off of practical prediction models. With such a noise model, the experimental method demonstrated could be used to quantify the economic advantage of forecast accuracy improvements to support decision making regarding data usage for the development of forecasting models in practical smart energy storage systems., and is not considered to be reflective of the behaviour of real prediction models.
@@ 4 Conclusions
This study investigated the impacts of data usage on the prediction accuracy of forecasting models for building operational conditions, and the resulting operational performance of a Model Predictive Control (MPC) scheme in a multi-building energy system with distributed generation and storage. Experiments were conducted using a large-scale dataset of electrical load measurements from buildings in the Cambridge University Estates.
@@
The relationship between forecast accuracy and operational performance of the MPC controller was investigated for synthetic forecasts, which showed the MPC scheme is most sensitive to grid electricity price and carbon intensity prediction accuracy. This analysis methodology would allow decision makers to determine whether expenditure on data collection to produce forecasting models is economic for a practical building energy system, i.e. whether the costs of data usage are outweighed by the improvements in operational performance provided.
@@ Appendix A Multi-building energy system asset specification
[…]
Building
Battery power capacity (kW)
Battery energy capacity (kW)
Battery efficiency (%)
Solar PV power capacity (kWp)
0
531
4246
90
461
3
60
478
90
38
9
14
108
90
20
11
342
2736
90
41
12
98
780
90
89
15
203
1622
90
143
16
306
2448
90
120
25
275
2198
90
118
26
1001
8006
90
136
32
157
1255
90
166
38
58
461
90
46
44
1084
8674
90
663
45
1245
9962
90
437
48
622
4973
90
500
49
57
458
90
258
@@ Appendix B MPC planning horizon
The planning horizon used in a MPC scheme determines its ability to effectively arbitrage energy and meet the operational objectives of energy management. Previous studies show that a planning horizon of 24 hours is sufficient when controlling HVAC systems oldewurtel2012UseModelPredictive , and recommend 48 hours as a suitable value for systems with distributed solar and storage thieblemont2017PredictiveControlStrategies .
@@ C.1 Simple neural models
For the simple neural models, implementations from the Pytorch Lighting library falcon2019PyTorchLightning were used. The models use an input window of 168 time steps (input dimension) and output a forecast of 48 time steps (output dimension). All models were trained using the Adam optimizer kingma2017AdamMethodStochastic , with a batch size of 32 and a learning rate determined using the built-in learning rate tuner.
@@ C.2 State-of-the-art machine learning models
For the complex, state-of-the-art models, implementations from the Pytorch Forecasting library beitner2024Jdb78Pytorchforecasting were used. The models take in data from the previous 72 time steps along with known covariate data from the coming 48 time steps, and output forecasts of the target variable for the coming 48 time steps. All models were trained using an early stopping trainer with a patience of 5, for a maximum of 50 epochs, with a limit of 800 training batches per epoch. The learning rate was determined using the built-in learning rate tuner. The size parameters, loss functions, and optimizers used are detailed in Table 4.
Model
TFT
NHiTS
DeepAR
Hidden size
48
128
64
Drop-out
0.1
0.1
0.1
Loss
Quantile loss
Quantile loss
Normal distribution loss
Optimizer
Adam
AdamW
Adam
Attention head size: 4
Weight decay: 0.01 Backcast loss ratio: 0
RNN layers: 3
@@ D.1 Functional data analysis
The fPCA approach used involves first aligning the data samples to a common mean, $\mu(t)$. This generates a warping function and an amplitude function for each data sample that describe how the data sample maps to the mean function. The warping function describes the phase relationship, i.e. the variation in time, and the amplitude function describes changes in magnitude. The warping and amplitude functions are then analysed separately and fPCs generated for both. The approach is illustrated schematically in Fig. 12, and full details of the approach are described in (Ward, 2021) ward2021DatacentricStochasticModel .
@@ D.2 Wasserstein similarity metric
The deconstruction of the data into fPCs and corresponding coefficients enables the use of standard statistical techniques to define measures of similarity between load profiles in the dataset. A load profile similarity metric based on an optimal transport approach is proposed. This metric is based on the approximate Wasserstein distance, or Earth-mover’s distance, between the distributions of PC coefficients, which provides a measure of the ease with which one probability distribution can be transformed into the other, and hence is a measure of the similarity of the two distributions. As such, the metric is referred to as the ‘Wasserstein similarity metric’, and is computed using the Geomloss feydy2024JeanfeydyGeomloss library. Advantages of this method are that it provides a comparison of the underlying patterns in the load profiles that is not excessively distorted by sequence misalignments as is the case for simple metrics like RMSE, and it allows time series of different durations to be compared.
@@ Appendix E Change-point analysis
A change-point analysis is conducted on the building electrical load dataset to detect changes in the load patterns in each building. The Bayesian ensemble algorithm “Bayesian Estimator of Abrupt change, Seasonal change, and Trend” (BEAST) zhao2019DetectingChangepointTrend is used to decompose the load time series and derive non-linear dynamics across multiple timescales, detecting change-points, seasonality, and trends. The value of this ensemble approach is that it quantifies the relative usefulness of individual decomposition models, leveraging all the models via Bayesian model averaging. Additionally, it provides an estimated probability of inferred change-points being ‘true’ change-points. Results of the change-point analysis for Buildings 26 & 44 are shown in Fig. 18. The demand profiles are decomposed into seasonal and trend components, and detected change-points for the two components are indicated in red and green respectively. The trend component of Building 44 is found to vary substantially more than that of Building 26, which corresponds with the visual variability of the demand profiles.

### v2

@@ 3.3.3 Online training
During building operation monitoring systems continuously collect operational data. The characteristics of building behaviour can change during operation due to external factors such as weather & climate, occupancy, and equipment degradation & maintenance. Continuous online training updates the predictive models using the collected monitoring data, improving the model’s ability to adapt to dynamic changes in building behaviour. A version of the Linear prediction model with online training is implemented to investigate the impact of online training frequency on model accuracy. The model is updated online by tuning the model parameters (retraining) using a sliding window of training data with length equal to the update frequency. The prediction accuracy of models updated at different frequencies is plotted in Fig. 9. Higher update frequencies led to increased prediction accuracy across all prediction variables. For example, prediction accuracy for grid electricity price improved by 15.7%, 14.0%, 12.1%, and 8.1% when updated monthly, quarterly, semi-annually, and annually, respectively, compared to the baseline model which is not trained online. Online training allows the prediction model to adapt to changes in the underlying trends which occur over time, as it can learn the trends in recent data. This improves prediction accuracy as the model does not need to generalise in time, as it is trained (updated) on data that is representative of the current prediction horizon. Therefore, it is expected that the additional computational hardware and system complexity required for online trained prediction models will be worthwhile in practical systems.
@@ 3.4 Sensitivity of Model Predictive Control to forecast accuracy
Whether investments in additional data to improve forecast accuracy provide net benefit to the operation of a smart energy storage system depends on the resulting improvements in operational performance achieved by the MPC controller. The determination of optimal data collection strategies therefore requires quantification of the relationship between forecast accuracy and MPC operational performance. To investigate this in a controlled setting, the operational performance achieved by the tested MPC scheme is evaluated with the use of synthetic forecasts. The synthetic forecasts are produced by adding a Gaussian random walk noise component to the ground-truth values, as described in Eq. 4,
@@
The results show that the tested MPC scheme is most sensitive to the forecast accuracy of the grid electricity price and carbon intensity variables, suggesting that the most resource and expense should be invested in producing accurate forecasts of the grid conditions. Additionally, whilst the prediction models tested in this study performed worst when forecasting solar generation, see Fig. 2b, it may not be worth expending additional resource to improve these forecasts as the MPC scheme is least sensitive to this variable. Though the grid component of operational performance is found to be by far the most sensitive to forecast accuracy, this is likely due to the synthetic forecast model used [^3] In the synthetic forecast model used, the Gaussian random walk noise component is regenerated at each time instance. This leads to the generation of forecasts which are correct on average, but can cause the MPC controller to take opposing energy management strategies in successive time steps, leading to large magnitudes of energy flows from the battery to correct the strategy. Further work is required to create a noise model for synthetic forecasts that is calibrated to the operational performance vs forecast accuracy trade-off of practical prediction models. With such a noise model, the experimental method demonstrated could be used to quantify the economic advantage of forecast accuracy improvements to support decision making regarding data collection for the development of forecasting models in practical smart energy storage systems., and is not considered to be reflective of the behaviour of real prediction models.
@@ 4 Conclusions
This study investigated the impacts of data on the prediction accuracy of forecasting models for building operational conditions, and the resulting operational performance of a Model Predictive Control (MPC) scheme in a multi-building energy system with distributed generation and storage. Experiments were conducted using a large-scale dataset of electrical load measurements from buildings in the Cambridge University Estates.
@@
The relationship between forecast accuracy and operational performance of the MPC controller was investigated for synthetic forecasts, which showed the MPC scheme is most sensitive to grid electricity price and carbon intensity prediction accuracy. This analysis methodology would allow decision makers to determine whether expenditure on data collection to produce forecasting models is economic for a practical building energy system, i.e. whether the costs of data are outweighed by the improvements in operational performance provided.
@@ Appendix A Multi-building energy system asset specification
The specifications of the distributed generation and storage assets for the multi-building energy system simulated in this study are provided in Table 3. The battery power capacities are set to be 3 times the mean building electrical load in the test dataset, the battery energy capacities are set to be able to provide 24 hours of that mean electrical load, and the solar PV power capacities are set by assuming 90% of the roof area of each building is filled by solar panels of power density 0.15 kWp/m2.
Building Battery power capacity (kW) Battery energy capacity (kW) Battery efficiency (%) Solar PV power capacity (kWp)
0 531 4246 90 461
3 60 478 90 38
9 14 108 90 20
11 342 2736 90 41
12 98 780 90 89
15 203 1622 90 143
16 306 2448 90 120
25 275 2198 90 118
26 1001 8006 90 136
32 157 1255 90 166
38 58 461 90 46
44 1084 8674 90 663
45 1245 9962 90 437
48 622 4973 90 500
49 57 458 90 258
@@ Appendix B MPC planning horizon
The planning horizon used in a MPC scheme determines its ability to effectively arbitrage energy and meet the operational objectives of energy management. Previous studies show that a planning horizon of 24 hours is sufficient when controlling HVAC systems [17], and recommend 48 hours as a suitable value for systems with distributed solar and storage [20].
@@ C.1 Simple neural models
For the simple neural models, implementations from the Pytorch Lighting library [63] were used. The models use an input window of 168 time steps (input dimension) and output a forecast of 48 time steps (output dimension). All models were trained using the Adam optimizer [64], with a batch size of 32 and a learning rate determined using the built-in learning rate tuner.
@@ C.2 State-of-the-art machine learning models
For the complex, state-of-the-art models, implementations from the Pytorch Forecasting library [65] were used. The models take in data from the previous 72 time steps along with known covariate data from the coming 48 time steps, and output forecasts of the target variable for the coming 48 time steps. All models were trained using an early stopping trainer with a patience of 5, for a maximum of 50 epochs, with a limit of 800 training batches per epoch. The learning rate was determined using the built-in learning rate tuner. The size parameters, loss functions, and optimizers used are detailed in Table 4.
Model TFT NHiTS DeepAR
Hidden size 48 128 64
Drop-out 0.1 0.1 0.1
Loss Quantile loss Quantile loss Normal distribution loss
Optimizer Adam AdamW Adam
Attention head size: 4 Weight decay: 0.01 Backcast loss ratio: 0 RNN layers: 3
@@ D.1 Functional data analysis
The functional Principal Component Analysis (fPCA) approach used involves first aligning the data samples to a common mean, $\mu(t)$. This generates a warping function and an amplitude function for each data sample that describe how the data sample maps to the mean function. The warping function describes the phase relationship, i.e. the variation in time, and the amplitude function describes changes in magnitude. The warping and amplitude functions are then analysed separately and fPCs generated for both. The approach is illustrated schematically in Fig. 12, and full details of the approach are described in (Ward, 2021) [66].
@@ D.2 Wasserstein similarity metric
The deconstruction of the data into fPCs and corresponding coefficients enables the use of standard statistical techniques to define measures of similarity between load profiles in the dataset. A load profile similarity metric based on an optimal transport approach is proposed. This metric is based on the approximate Wasserstein distance, or Earth-mover’s distance, between the distributions of PC coefficients, which provides a measure of the ease with which one probability distribution can be transformed into the other, and hence is a measure of the similarity of the two distributions. As such, the metric is referred to as the ‘Wasserstein similarity metric’, and is computed using the Geomloss [67] library. Advantages of this method are that it provides a comparison of the underlying patterns in the load profiles that is not excessively distorted by sequence misalignments as is the case for simple metrics like RMSE, and it allows time series of different durations to be compared.
@@ Appendix E Change-point analysis
A change-point analysis is conducted on the building electrical load dataset to detect changes in the load patterns in each building. The Bayesian ensemble algorithm “Bayesian Estimator of Abrupt change, Seasonal change, and Trend” (BEAST) [57] is used to decompose the load time series and derive non-linear dynamics across multiple timescales, detecting change-points, seasonality, and trends. The value of this ensemble approach is that it quantifies the relative usefulness of individual decomposition models, leveraging all the models via Bayesian model averaging. Additionally, it provides an estimated probability of inferred change-points being ‘true’ change-points. Results of the change-point analysis for Buildings 26 & 44 are shown in Fig. 18. The demand profiles are decomposed into seasonal and trend components, and detected change-points for the two components are indicated in red and green respectively. The trend component of Building 44 is found to vary substantially more than that of Building 26, which corresponds with the visual variability of the demand profiles.